Amy has a pen 15 cm long that she wants to send as a gift to a friend. She finds a rectangular box in which the pen just fits. The length and width of the box are 5cm and 14cm respectively. What is the height of the box?
Answer by Boreal(15235)

It helps to draw this
The diagonal of the floor of the box is the sqrt (5^2+14^2)=sqrt (221)
That hypotenuse becomes the leg of a triangle whose other leg is the unknown height and the diagonal is the pen itself.
so sqrt(221)^2+h^2=15^2
221+h^2=225
h=2 cm.